OpenStudy (anonymous):

Solve each equation by completing the square. 8. x^2-6x=-4 9. x^2 +8 =6x 10. 2x^2-20x=8 11.x^2=24-4x 12. 10x + x^2 =42 13. 2x^2 + 8x -15 =0

OpenStudy (anonymous):

completing the square is adding or subtracting the constant to the other side, then using (b/2)^2 to make it factorable. Obviously doing all six of those is going to take too much time, so you should be able to do the rest. >.> x^2-6x=-4 add the (b/2)^2 x^2-6x+9=-4 factor it (x-3)(x-3)=-4 set each factor (in this case, both are x--3 so you only need to do it once because they're the same) equal to -4 x-3=-4 add +3 to both sides x=-1